The sum of two numbers is 2, and their difference is 17.
Answer by ewatrrr(24785)

 
Hi
let x and y represent the numbers
Question States***
x + y = 2
x - y = 17
2x = 19
x = 19/2 and y = -15/2 ( y = 4/2-19/2)
CHECKING our Answer***
1 + y = 2
19/2 + -15/2 = 4/2 = 2